OpenStudy (anonymous):

Find the absolute extrema of the function (x^2+4)^1/3 on the interval [-5,4]

OpenStudy (anonymous):

The absolute max and min.

OpenStudy (anonymous):

take the derivative, find the critical points, check those and also check the endpoints

OpenStudy (anonymous):

although there is really not that much to check this is an even function so it is smallest at \(x=0\) by pure reason, no calculus

OpenStudy (anonymous):

The derivative I got is \[1/3(X^2+4)^-2/3 (2X)\] I'm not sure how to solve to get the critical points.

OpenStudy (anonymous):

you have \(f(x)=\sqrt[3]{x^2+4}\) and it is pretty obvious that this is largest when \(x^2+4\) is largest, and smallest when \(x^2+4\) is smallest

OpenStudy (anonymous):

\(x^2+4\) will be smallest when \(x=0\) and largest when \(x=-5\) we can do it the stupid calculus way if you like

OpenStudy (anonymous):

you have the derivative correct (ugly but correct)

OpenStudy (anonymous):

use fraction notation \[f'(x)=\frac{2x}{\sqrt[3]{(x^2+4)^2}}\]

OpenStudy (anonymous):

Lol ah okay that makes sense so then would there only be one critical point? x=0?

OpenStudy (anonymous):

now the critical points are where it is undefined (where the denominator is zero) or where it is zero (where the numerator is zero

OpenStudy (anonymous):

right, only at \(x=0\)

OpenStudy (anonymous):

since \(x^2+4\geq 4\) the denominator is never zero

OpenStudy (anonymous):

So now I have to plug in 0 and the end points into the original equation right.

OpenStudy (anonymous):

but i think if you were not in calculus class, the answer would be totally obvious the cubed root of something is largest when that something is largest right? and on that interval \(x^2+4\) is largest if \(x=-5\)

OpenStudy (anonymous):

yes plug it in to the ORIGINAL function not the derivative

OpenStudy (anonymous):

Is there any way I can check myself on my TI-84 calculator?

OpenStudy (anonymous):

if you plug the critical point in to the derivative you get zero since thats how you got it in the first place

OpenStudy (anonymous):

calculators are so last century http://www.wolframalpha.com/input/?i=%28x^2%2B4%29^1%2F3

OpenStudy (anonymous):

this is even better for your interval http://www.wolframalpha.com/input/?i=%28x^2%2B4%29^1%2F3+domain+-5..4

OpenStudy (anonymous):

Well the only reason I'm even asking is so I can check myself on the test and because every time I plug all the numbers into the original equation and get the answers and then try use my calculator to solve also I get different answers. So I thought maybe that I was plugging it into my calculator wrong. The min values I got are (0,1.6) and (4,1.6)